# Gamma-ray burst pulse structures and emission mechanisms

### v1

@@ 1 Introduction
Gamma-ray bursts (GRBs) are well known to have erratic light curves in the X- and $\gamma$-ray bands during their prompt emission phase. While the light curves of some GRBs are smooth (i.e., contain a single pulse), the light curves of many others contain multiple, sometimes overlapping pulses. These pulses are considered to be individual emission episodes varying largely in duration and energy. This makes it very hard to deduce the physical properties of GRBs, as a complete model requires understanding the origin of individual pulse duration, shape, spectrum and the relation between the different pulses.
Studies of individual pulses have proven fruitful in revealing some key properties. In early works using data from the BATSE instrument on the Compton Gamma-Ray Observatory, Norris et al. (1996, 2005) and Hakkila et al. (2018) found that individual pulses are typically characterized by non-symmetric shape, which is modelled as ‘fast rise exponential decay’ (FRED). This shape was found independent of the observed energy band (Hakkila & Cumbee, 2009). A FRED pulse shape can result from a sudden dissipation of energy. This idea indeed suggests a non-symmetric pulse shape: While the rise in the light curve is due to the unspecified energy dissipation mechanism, the decay can be affected by additional physical effects, such as light aberration (known as high-latitude emission; e.g., Panaitescu & Kumar, 2001).
@@
Additional observational findings that may shed light on the physical origin of the pulses include: (i) pulses detected at lower energies tend to have longer duration compared to pulses observed at higher energies (Richardson et al., 1996). (ii) A correlation was found between pulse fluence and duration (Hakkila & Preece, 2011): brighter pulses tend to be longer. And (iii) Pulse duration were found to be anti-correlated with their peak fluxes (Hakkila & Preece, 2011).
Along with these temporal analyses of the GRB pulses light curves, Li et al. (2021) employed a time-resolved spectral analysis method in analyzing multi-pulsed GRBs observed by Fermi-GBM. They found a clear trend between the low energy spectral index $\alpha$ and the pulse number (first, second, third, etc.). As a burst evolves and the pulse number increases, the value of $\alpha$ decreases. This indicates a temporal evolution in the radiative mechanism across pulses during the prompt phase in a single burst.
Most of the above-mentioned works dealt with analyzing data obtained from the BATSE instrument. However, some of the well-used models provide poor fits to data (Hintze, 2022) obtained by the much newer Gamma-ray Bursts Monitor (GBM) on board the Fermi satellite (Meegan et al., 2009). Indeed, it seems that very few studies have been published analyzing the shape of pulses observed by the GBM instrument. Clearly, being a newer instrument, the signal-to-noise (S/N) ratio of the data provided by the GBM is higher than that of the BATSE. This motivates a study of the GRB pulse shapes detected by the GBM, aiming to propose a more physically grounded model for fitting these pulse shapes and to explore various potential correlations between pulse shape, pulse count, duration, and spectral characteristics. Focusing on multi-pulse evolution enables one to study if the changes in pulse structures correspond to changes in emission mechanism within the relativistic flow.
@@
By definition, the first photons are always observed from the photosphere, whose radius is expected in the range $r_{ph}\sim 10^{11}-10^{13}$ cm, depending on the outflow parameters (e.g., Pe’er et al., 2012). The expected spectrum is a modified black-body. A “pure” black-body (Planck) function is not expected due to both light aberration (Pe’er, 2008) and possible energy dissipation below and close to the photosphere (Pe’er et al., 2006; Giannios, 2006). Energy dissipation at larger radii, $r_{\gamma}\approx 10^{13}-10^{17}$ cm will occur in the optically thin region and be observed after a few to several hundred seconds. Being in the optically thin regime, the resulting spectra reflect the radiative process and are expected to be a combination of synchrotron at lower energies and IC at higher energies. In principle, in extended jetted outflows as is expected in long GRBs, the two emission models can coexist and contribute simultaneously to the observed spectra (e.g., Ryde, 2005; Ryde & Pe’er, 2009; Guiriec et al., 2011; Ajello et al., 2020).
The value of the low-energy spectral index ($\alpha$), which is typically well measured, can be used as a proxy for the emission mechanism: if the emission is of thermal (photospheric) origin, $\alpha$ represents the Rayleigh-Jeans part of the spectrum (although modified) and is therefore expected to be harder than the expected index from synchrotron emission. We point out that various effects, both physical, geometrical and instrumental, will act to modify the observed value of $\alpha$, and therefore, the discrimination between the radiative models is done using empirical models fitted to the synthetic data. The boundaries of synchrotron emission would lie below $\alpha\leq-0.66$ (known as the “synchrotron line-of-death”). On the other hand, it was shown by Acuner et al. (2020) that $\alpha\geq-0.1$ is a clear indication of a non-dissipative photospheric (NDP) origin. This value is modified from the theoretical limit of $\alpha=+0.4$ (Beloborodov, 2011) constrained by the finite detector’s energy band, the spectral fitting procedure (they used a cut-off power law), and the finite S/N ratio. In order not to be sensitive to these unknowns, we interpret $\alpha$ of indicating thermal origin using only the non-modified theoretical limit, $\alpha\geq+0.4$; see further discussion below.
@@ 2.1 Sample selection
We use data obtained by the Fermi Gamma-Ray Burst Monitor (GBM), which includes 12 NaI and 2 BGO detectors covering roughly the energy range between 8 keV to 40 MeV (Meegan et al., 2009). We define the basic unit of a burst morphology as a ‘pulse’, which is considered as an individual emission episode. Ideally, a pulse’s count-rate should start to rise from the background level, reach a peak and then decay to background levels. Therefore, most of the pulses have overlapping emission activities that can cause subdominant variations. The pulse duration is defined as the time interval from the peak count rate to half of the peak count rate on both the ascending (rise) and descending (decay) portions of the pulse. We denote the duration of a pulse as the source interval.
In our analysis, we define the pulses as follows: the pulse count-rate should rise from a background level, if not, then immediately following another pulse. In this second case, the rise phase is identified if, at the start time of the pulse, it contains at most 50% of the peak count-rate. Similarly, the count-rate at the end of the decay phase should be similar to the background levels. If not (namely, there is a following overlapping pulse), then the end time of the pulse should be at most 50% of the peak count-rate immediately followed by another pulse  [^1] Typically, at the start of the rise phase and end of the decay phase, the count-rates are much lower than these maximal values.. To mitigate the effects of light curve binning and background noisy flares in the initial data, a 2% error is allowed in this criterion. If one of these criteria is not met, we do not claim an identification of an individual pulse. These criteria are set to maximize the duration of individual pulses while maintaining a separation between partially overlapping pulses.
Furthermore, these criteria effectively act as a low pass filter, enabling smooth fluctuations in the light curve. As suggested by Li et al. (2021), we are allowing pulses with smaller spikes whose heights are bounded by an approximate pulse-shaped envelope, identified after the 0.1s binning of the brightest NaI detector light curve. This avoids statistical variations that could be produced in the light curve due to background and ensures a high quality of individual pulses in a burst, for this analysis.
@@
After identifying the pulses, we impose additional selection criteria as follows. The light curves are binned using Bayesian blocks (Scargle et al., 2013); see further discussion in section 2.3 below. A minimum selection criterion is that each pulse in each GRB light curve has at least two time bins, each of them having at least a statistical significance $S\geq 20$. An exact definition of $S$, which is a measure of the signal-to-noise ratio, can be found in Vianello (2018). Secondly, we selected only pulses which have the coefficient of determination $r^{2}$ value $\geq$ 0.7 222 The coefficient of determination, denoted by $r^{2}$ $(0\leq r^{2}\leq 1)$ is a measure of how well a function fits a given data. It is defined by $r^{2}\equiv 1-{RSS}/{TSS}$, where $RSS$ is the sum of squares of the fit residuals, and $TSS$ is the total sum of squares. In our analysis, the pulses are selected only if $r^{2}>0.7$. after fitting with our novel pulse shape function. We also removed the fits with significant errors. This criterion ensures that the individual pulses are easily distinguishable in the burst and enables us to carry out a relatively “clean” analysis, i.e., that the errors introduced in the fitting procedure are slight.
@@ 2.2 Pulse Model Analysis
In order to analyse the pulses detected by the GBM instrument, we introduce a new fitting function different from the one used previously in the literature. Our new function is composed of two approximate sigmoid functions and is given in Equation 1. It can properly fit for the slopes, width and height of a pulse. Its main advantage over the fit function used in the literature (Norris et al., 1996) is that it properly captures wide pulse shapes, which is reflected in a more accurate description of the rising and decaying slopes. It is capable of fitting for both FRED and symmetric pulse shapes as the parameters are flexible.
@@
$I(t)=\frac{A}{4}\times\left[{1-tanh\left(\frac{1}{s_{r}}(t-r_{r})\right)}\right]\times\left[{1+tanh\left(\frac{1}{s_{l}}(t-r_{l})\right)}\right].$ (1)
Here, I(t) is given in units of count-rate, $s_{l}$ and $s_{r}$ represent the rising (pre-peak) and decaying (post-peak) slopes of the light curve - smaller values imply steeper slopes (before and after the peak, respectively). The parameters $r_{l}$ and $r_{r}$ are the half-time of light curve rise and fall, respectively, and $A$ is a global normalization. The subscripts $l$ and $r$ indicate the left and right sides of the peak respectively. This function, therefore, has a total of five degrees of freedom. It is designed for best fits of a single peak pulse, enabling it to capture both symmetric and FRED-like behaviour on an equal footing. Examples of pulse shapes that can be obtained from fitting the function are given in Figure 2. Furthermore, this function enables good fits to pulses with elongated peaks (see examples in Figure 3), often manifested by multiple light curve fluctuations.
@@
Figure 3: Examples of fitting different pulses from the sample according to the pulse shape function. The solid green line represents the raw count-rate data from GBM, and the dashed blue line represents the pulse shape model fit according to Equation 1. Left: The novel pulse function fit for the prompt light curve of GRB 081009 (bn 08109140). Here, the pulse shape parameter $\phi$ = 0.35, the pulse shape is FRED-like and it is a single-peaked pulse. Middle: The novel pulse function fit for the prompt light curve of GRB 190114 (bn 190114873). Here, the pulse shape parameter $\phi$ = 2.12, and the pulse shape is Symmetric-like. Right: The novel pulse function fits the prompt light curve of GRB 120328 (bn 120328268). The pulse shape parameter $\phi$ = 0.17, and the pulse shape is combined-peak pulse.
@@
We assume that the subdominant peaks with a height that does not exceed 20% of the peak count-rate of the pulse in 0.1s binning of the light curve are part of the dominant pulse emission. Any pulse with fluctuations that rise and decay in less than a second and with peak heights that exceed this 20% level of peak count-rate is categorized as a combined peak pulse. These values are based on the background count-rate before and after the pulse duration. This categorisation of pulses enables us to be aware of the variations that we may have smoothed over while defining a pulse. Some pulses could have peak heights marginally over the bound and cannot be categorised into either type. For example, in Figure 3, the first subfigure of GRB 081009(140) is a single peak pulse, while the third one of GRB 130606(497) is a combined peak pulse. The second example of GRB 190114(873) represents a pulse which cannot be categorised into either type. The burst name, pulse number, start and end times of each pulse, and the corresponding fitting parameters A, $s_{l}$, $s_{r}$, $r_{l}$, $r_{r}$, the goodness of fit measurement $r^{2}$ after the fit to Equation 1, along with the shape $\phi$ are given in Table 1.
@@ 3.2 The pulse shape studies
Figure 6: Left: $s_{l}$ (rising slope) - Pulse number relation. The green and orange dots represent single and combined peak pulses respectively. The last pulses of each GRBs are also marked here. Right: $s_{r}$ (decaying slope)- Pulse number relation. The green and orange dots represent single and combined peak pulses respectively. The last pulses of each GRBs are also marked here.

### v2

@@ 1 Introduction
Gamma-ray bursts (GRBs) are well known to have erratic light curves in the X- and $\gamma$-ray bands during their prompt emission phase. While the light curves of some GRBs are smooth (i.e., contain a single pulse), the light curves of many others contain multiple, sometimes overlapping pulses. These pulses are considered to be individual emission episodes varying largely in duration and energy. This makes it very hard to deduce the physical properties of GRBs, as a complete model requires understanding the origin of individual pulse duration, shape, spectrum, and the relation between the different pulses.
Studies of individual pulses have proven fruitful in revealing some key properties. In early works using data from the BATSE instrument on the Compton Gamma-Ray Observatory, Norris et al. (1996, 2005) and Hakkila et al. (2018) found that individual pulses are typically characterized by a non-symmetric shape, which is modelled as ‘fast rise exponential decay’ (FRED). This shape was found independent of the observed energy band (Hakkila & Cumbee, 2009). A FRED pulse shape can result from a sudden dissipation of energy. This idea indeed suggests a non-symmetric pulse shape: While the rise in the light curve is due to the unspecified energy dissipation mechanism, the decay can be affected by additional physical effects, such as light aberration (known as high-latitude emission; e.g., Panaitescu & Kumar, 2001).
@@
Additional observational findings that may shed light on the physical origin of the pulses include: (i) Pulses detected at lower energies tend to have longer duration compared to pulses observed at higher energies (Richardson et al., 1996). (ii) A correlation was found between pulse fluence and duration (Hakkila & Preece, 2011): brighter pulses tend to be longer. And (iii) Pulse durations were found to be anti-correlated with their peak fluxes (Hakkila & Preece, 2011).
Along with these temporal analyses of the GRB pulses’ light curves, Li et al. (2021) employed a time-resolved spectral analysis method in analyzing multi-pulsed GRBs observed by Fermi-GBM. They found a clear trend between the low energy spectral index $\alpha$ and the pulse number (first, second, third, etc.). As a burst evolves and the pulse number increases, the value of $\alpha$ decreases. This indicates a temporal evolution in the radiative mechanism across pulses during the prompt phase in a single burst.
Most of the above-mentioned works dealt with analyzing data obtained from the BATSE instrument. However, some of the well-used models provide poor fits to data (Hintze, 2022) obtained by the much newer Gamma-ray Bursts Monitor (GBM) on board the Fermi satellite (Meegan et al., 2009). Indeed, it seems that very few studies have been published analyzing the shape of pulses observed by the GBM instrument. Clearly, being a newer instrument, the signal-to-noise (S/N) ratio of the data provided by the GBM is higher than that of the BATSE. This motivates a study of the GRB pulse shapes detected by the GBM, aiming to propose a more physically grounded model for fitting these pulse shapes and to explore various potential correlations between pulse shape, pulse count, duration, and spectral characteristics. Focusing on multi-pulse evolution enables one to study whether the changes in pulse structures correspond to changes in the emission mechanism within the relativistic flow.
@@
By definition, the first photons are always observed from the photosphere, whose radius is expected in the range $r_{ph}\sim 10^{11}-10^{13}$ cm, depending on the outflow parameters (e.g., Pe’er et al., 2012). The expected spectrum is a modified black-body. A “pure” black-body (Planck) function is not expected due to both light aberration (Pe’er, 2008) and possible energy dissipation below and close to the photosphere (Pe’er et al., 2006; Giannios, 2006). Energy dissipation at larger radii, $r_{\gamma}\approx 10^{13}-10^{17}$ cm, will occur in the optically thin region and be observed after a few to several hundred seconds. Being in the optically thin regime, the resulting spectra reflect the radiative process and are expected to be a combination of synchrotron at lower energies and IC at higher energies. In principle, in extended jetted outflows as is expected in long GRBs, the two emission models can coexist and contribute simultaneously to the observed spectra (e.g., Ryde, 2005; Ryde & Pe’er, 2009; Guiriec et al., 2011; Ajello et al., 2020).
The value of the low-energy spectral index ($\alpha$), which is typically well measured, can be used as a proxy for the emission mechanism: if the emission is of thermal (photospheric) origin, $\alpha$ represents the Rayleigh-Jeans part of the spectrum (although modified) and is therefore expected to be harder than the expected index from synchrotron emission. We point out that various effects, both physical, geometrical, and instrumental, will act to modify the observed value of $\alpha$, and therefore, the discrimination between the radiative models is done using empirical models fitted to the synthetic data. The boundaries of synchrotron emission would lie below $\alpha\leq-0.66$ (known as the “synchrotron line-of-death”). On the other hand, it was shown by Acuner et al. (2020) that $\alpha\geq-0.1$ is a clear indication of a non-dissipative photospheric (NDP) origin. This value is modified from the theoretical limit of $\alpha=+0.4$ (Beloborodov, 2011) constrained by the finite detector’s energy band, the spectral fitting procedure (they used a cut-off power law), and the finite S/N ratio. In order not to be sensitive to these unknowns, we interpret $\alpha$ of indicating thermal origin using only the non-modified theoretical limit, $\alpha\geq+0.4$; see further discussion below.
@@ 2.1 Sample selection
We use data obtained by the Fermi Gamma-Ray Burst Monitor (GBM), which includes 12 NaI and 2 BGO detectors covering roughly the energy range between 8 keV to 40 MeV (Meegan et al., 2009). We define the basic unit of a burst morphology as a ‘pulse’, which is considered as an individual emission episode. This is part of the light curve that rises, reaches a peak, and decays. Due to the variability in the pulse structures in GRBs, the method for pulse demarcations has been mostly visual in the previous studies. As the reproducibility of pulses is of prominent importance, we have made specific criteria for the count rate to ensure that the general pulse envelope can be identified.
  • Ideally, a pulse is defined as a distinct emission episode characterized by a rise and fall in the count-rate relative to the background level. The pulse start time is defined as the time at which the pulse count-rate sustainably rises above the background level. The effects of Poissonian noise in pulse selection are taken care of by the application of a polynomial fitting to the background level and the Bayesian block binning method. This enables us to identify the beginning of the pulse. The pulse count rate rises until it reaches a peak. Following, the pulse count rate descends, marking the beginning of the decay phase. The pulse end time is set once the count rate reaches again the background level.
  • In some cases, the pulse count-rate does not start or end at background levels. This happens when one pulse immediately follows another pulse. In order to identify whether an emission episode forms an individual, independent pulse or should be considered as a fluctuation within an already identified pulse, the following criterion is applied. We measure the count rates at the beginning and end of the emission episode.
    • To consider the emission episode that immediately follows a pulse as another, independent pulse, we calculate the count rates at the start time of the episode and at its peak. Our criterion is that the count rate at the start time must be less than 50% of the peak count rate. If this criterion is met, we consider the episode as an independent pulse. Otherwise, we consider it as a continuation of the previous pulse or as a subdominant variation.
    • Similarly, the emission episode is considered a separated pulse if at the end time of the episode, the count rate is at most 50% of the peak count-rate.
To mitigate the effects of light-curve binning and background noisy flares in the initial data, a 2% error is allowed in this criterion, see Section 2.2. We denote the duration of a pulse as the source interval.
Furthermore, these criteria effectively act as a low-pass filter, enabling smooth fluctuations in the light curve. As suggested by Li et al. (2021), we are allowing pulses with smaller spikes whose heights are bounded by an approximate pulse-shaped envelope, identified after the 0.1s binning of the brightest NaI detector light curve. This avoids statistical variations that could be produced in the light curve due to background and ensures a high quality of individual pulses in a burst, for this analysis.
@@
After identifying the pulses, we impose additional selection criteria as follows. The light curves are binned using Bayesian blocks (Scargle et al., 2013); see further discussion in section 2.3 below. A minimum selection criterion is that each pulse in each GRB light curve has at least two time bins, each of them having at least a statistical significance $S\geq 20$. An exact definition of $S$, which is a measure of the signal-to-noise ratio, can be found in Vianello (2018). Secondly, we selected only pulses which have the coefficient of determination $r^{2}$ value $\geq$ 0.7 111 The coefficient of determination, denoted by $r^{2}$ $(0\leq r^{2}\leq 1)$ is a measure of how well a function fits a given data. It is defined by $r^{2}\equiv 1-{RSS}/{TSS}$, where $RSS$ is the sum of squares of the fit residuals, and $TSS$ is the total sum of squares. In our analysis, the pulses are selected only if $r^{2}>0.7$. after fitting with our novel pulse shape function. We also removed the fits with significant errors. This criterion ensures that the individual pulses are easily distinguishable in the burst and enables us to carry out a relatively “clean” analysis, i.e., that the errors introduced in the fitting procedure are slight.
@@ 2.2 Pulse Model Analysis
In order to analyse the pulses detected by the GBM instrument, we introduce a new fitting function different from the one used previously in the literature. Our new function is composed of two approximate sigmoid functions and is given in Equation 1. It can properly fit for the slopes, width and height of a pulse. It is capable of fitting for both FRED and symmetric pulse shapes as the parameters are flexible.
@@
$I(t)=\frac{A}{4}\times\left[{1-\tanh\left(\frac{1}{s_{r}}(t-r_{r})\right)}\right]\times\left[{1+\tanh\left(\frac{1}{s_{l}}(t-r_{l})\right)}\right].$ (1)
Here, I(t) is given in units of count-rate, $s_{l}$ and $s_{r}$ represent the rising (pre-peak) and decaying (post-peak) slopes of the light curve - smaller values imply steeper slopes (before and after the peak, respectively). The parameters $r_{l}$ and $r_{r}$ are the half-time of light curve rise and fall, respectively, and $A$ is a global normalization. The subscripts $l$ and $r$ indicate the left and right sides of the peak, respectively. This function, therefore, has a total of five degrees of freedom. It is designed for the best fits of a single peak pulse, enabling it to capture both symmetric and FRED-like behaviour on an equal footing. Examples of pulse shapes that can be obtained from fitting the function are given in Figure 2. Furthermore, this function enables good fits to pulses with elongated peaks (see examples in Figure 3), often manifested by multiple light curve fluctuations.
@@
Figure 3: Examples of fitting different pulses from the sample according to the pulse shape function. The solid green line represents the raw count-rate data from GBM, and the dashed blue line represents the pulse shape model fit according to Equation 1. Left: The novel pulse function fit for the prompt light curve of GRB 081009 (bn 08109140). Here, the pulse shape parameter $\phi$ = 0.35, the pulse shape is FRED-like and it is a single-peaked pulse. Middle: The novel pulse function fit for the prompt light curve of GRB 190114 (bn 190114873). Here, the pulse shape parameter $\phi$ = 2.12, and the pulse shape is Symmetric-like. Right: The novel pulse function fits the prompt light curve of GRB 120328 (bn 120328268). The pulse shape parameter $\phi$ = 0.17, and the pulse shape is FRED-like and it is a combined-peak pulse.
@@
We assume that the subdominant peaks with a height that does not exceed 20% of the peak count-rate of the pulse in 0.1s binning of the light curve are part of the dominant pulse emission. Any pulse with fluctuations that rise and decay in less than a second and with peak heights that exceed this 20% level of peak count-rate is categorized as a combined peak pulse. This categorisation of pulses enables us to be aware of the variations that we may have smoothed over while defining a pulse. Some pulses could have peak heights marginally over the bound and cannot be categorised into either type. For example, in Figure 3, the first subfigure of GRB 081009(140) is a single peak pulse, while the third one of GRB 130606(497) is a combined peak pulse. The second example of GRB 190114(873) represents a pulse which cannot be categorised into either type. We ensure that the signal-to-noise ratio is high by the criterion that at least one Bayesian bin of the pulse has $S>20$. This condition, however, does not guarantee that all the bins, especially the ones at the beginning and at the end of each pulse, have the required S/N ratio. The allowed range of the count-rate at the beginning and at the end of the pulse, 48 - 52 % ($50\pm 2~{}\%$) and the combined peak pulse classification takes into account these smaller subdominant variations and structures in the pulse envelope in our analysis. We have checked that this variation does not affect the obtained value of the rising or decaying slopes. Therefore, it does not affect our shape parameter, $\phi$.
Equation 1, while similar to a logistic function, offers more flexibility than a simple Gaussian or logistic function, by allowing both the rise and decay slopes to be varied independently. This property also allows one to fit the pulses which have extended or plateau-like peaks, such as Figure 3 [Right]. Although Equation 1 is a combination of two logistic functions, the resulting pulse shape $\phi$ parameter is more direct and more sensitive to the varied pulse shapes of GRBs than the traditional statistical measures like skewness or kurtosis. Moreover, Equation 1 has two main advantages over previous fit functions used in the literature (e.g., Norris et al., 1996, 2005). The function used by Norris et al. (2005) is an asymmetric function. As such, this function can not capture a pulse shape where the rise time exceeds the decay time, whereas Equation (1) accommodates such cases with a pulse shape parameter $\phi>1$. Furthermore, the fitting function used by Norris et al. (2005) characterizes the pulse asymmetries using a parameter $k$, where $k=0$ represents symmetric pulses and values in the range $10<k<$ 1 correspond to asymmetric pulses with decay times longer than rise times. The limited range of $k$ values makes distinguishing between symmetric-like and FRED-like pulses more challenging. In contrast, in the model used here, distinguishing between a FRED-like pulse shape ($0<\phi\leq 0.3$) and symmetric-like pulses ($\phi>1$) is more direct. We provide further details of the difference between the fitting functions in Appendix A.
The burst name, pulse number, start and end times of each pulse, and the corresponding fitting parameters A, $s_{l}$, $s_{r}$, $r_{l}$, $r_{r}$, the goodness of fit measurement $r^{2}$ after the fit to Equation 1, along with the shape $\phi$ are given in Table 1.
@@ 3.2 The pulse shape studies
Figure 6: Left: $s_{l}$ (rising slope) - Pulse number relation. The green and orange dots represent single and combined peak pulses, respectively. The last pulses of each GRBs are also marked here. Right: $s_{r}$ (decaying slope)- Pulse number relation. The green and orange dots represent single and combined peak pulses respectively. The last pulses of each GRBs are also marked here.